A simple random sample from a population with a normal distribution of 106 body temperatures has x=98.20°F and s=0.690 °F.

Construct a 99​% confidence interval estimate of the standard deviation of body temperature of all healthy humans.
Answer #1

Here s = 0.69 and n = 106
df = 106 - 1 = 105

α = 1 - 0.99 = 0.01
The critical values for α = 0.01 and df = 105 are ꭓ^2(1-α/2,n-1) = 71.4282 and ꭓ^2(/2,n-1) = 146.0696

(n − 1)sos (n - 1)s V xázs xá/2 xias

CI = (sqrt(105*0.69^2/146.0696) , sqrt(105*0.69^2/71.4282))
CI = (0.5850 , 0.8366)